Calculate P(0.2<=X<=0.4) for the following discrete probability mass function: X = {0.1, 0.2, 0.3, 0.4} and f(x) = {0.29, 0.15, 0.35, 0.21}
Answer #1

f (X) = P (X=x)

Given P (X = 0.1) = 0.29

P (X= 0.2) = 0.15

P (X= 0.3) = 0.35

P (X= 0.4) = 0.21

We need to find

P ( 0.2 0.4 ) = P (X=0.2)+P(X=0.3)+P (X=0.4)

= 0.15 + 0.35 + 0.21

= 0.71